### v6

@@ 1 Introduction
To address this challenge, we propose a logic-driven data augmentation method based on Abstract Meaning Representation (AMR). AMR is a structural representation of the semantics and logical structure of text via a rooted directed acyclic graph (DAG) Shou et al. (2022). Figure 1 shows an example of an AMR graph. The AMR graph can be easily modified by changing nodes or arguments to create logically equivalent or nonequivalent graphs. By taking advantage of the ease of logical manipulation of AMR graph and of end-to-end conversion between natural language and AMR graphs, our proposed data augmentation is not task-specific or template-dependent, and can generate logically equivalent and nonequivalent sentences that are diverse in their use of language.
@@ 2 Related Work
Logical reasoning is rigorous thinking to derive a conclusion based on a given premise Seel (2011); Bronkhorst et al. (2020). Existing reasoning datasets reasoning can be categorised into two levels: sentence level, including tasks like natural language inference that assess if one sentence logically follows from another (e.g., MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), QQP Wang et al. (2018)); passage level, which requires logical deduction from given contexts, questions, and multiple choices (e.g., PARARULE Clark et al. (2021), PARARULE-Plus Bao et al. (2022)) and reading comprehension tasks (e.g., ReClor Yu et al. (2020), LogiQA Liu et al. (2021)). We introduce an abstract meaning representation-based methodology for logic-driven data augmentation aimed at enhancing models’ logical reasoning capabilities across these tasks.
@@ 4.1 Datasets
ReClor Yu et al. (2020) and LogiQA Liu et al. (2021) are two challenging logical reasoning datasets. ReClor is collected from the Graduate Management Admission Test (GMAT) and Law School Admission Test (LSAT). LogiQA is collected from the National Civil Service Examination Liu et al. (2021). Additionally, we performed evaluation on five datasets for natural language inference and textual entailment tasks: MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), and QQP Wang et al. (2018). MNLI, RTE, and MRPC assess the relationship between two sentences, while QNLI focuses on the relationship between a question and a sentence, and QQP evaluates the relationship between two questions.
@@ 4.2 Settings
4.3 Logical-Equivalence-Identification Contrastive Learning for Discriminative LLM
@@ 4.4 Prompt Augmentation for Generative LLM
We assessed the robustness of AMR-LDA and LReasoner-LDA models on the PARARULE-Plus dataset Bao et al. (2022) by modifying the test set with the contraposition law. Examples from this dataset can be found in Appendix Figures 9 and 10. AMR-LDA showed enhanced robustness on these altered tests compared to LReasoner-LDA.

### v7

@@ 1 Introduction
To address this challenge, we propose a logic-driven data augmentation method based on Abstract Meaning Representation (AMR). AMR is a structural representation of the semantics and logical structure of text via a rooted directed acyclic graph (DAG) Shou et al. (2022). Figure 1 shows an example of an AMR graph. The AMR graph can be easily modified by changing nodes or arguments to create logically equivalent or nonequivalent graphs. By taking advantage of the ease of logical manipulation of AMR graphs and of end-to-end conversion between natural language and AMR graphs, our proposed data augmentation is not task-specific or template-dependent, and can generate logically equivalent and nonequivalent sentences that are diverse in their use of language.
@@ 2 Related Work
Logical reasoning is rigorous thinking to derive a conclusion based on a given premise Seel (2011); Bronkhorst et al. (2020). Existing reasoning datasets’ reasoning can be categorised into two levels: sentence level, including tasks like natural language inference that assess if one sentence logically follows from another (e.g., MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), QQP Wang et al. (2018)); passage level, which requires logical deduction from given contexts, questions, and multiple choices (e.g., PARARULE Clark et al. (2021), PARARULE-Plus Bao et al. (2022a)) and reading comprehension tasks (e.g., ReClor Yu et al. (2020), LogiQA Liu et al. (2021)). We introduce an abstract meaning representation-based methodology for logic-driven data augmentation aimed at enhancing models’ logical reasoning capabilities across these tasks.
@@ 4.1 Datasets
ReClor Yu et al. (2020) and LogiQA Liu et al. (2021) are two challenging logical reasoning datasets. ReClor is collected from the Graduate Management Admission Test (GMAT) and the Law School Admission Test (LSAT). LogiQA is collected from the National Civil Service Examination Liu et al. (2021). Additionally, we performed evaluations on five datasets for natural language inference and textual entailment tasks: MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), and QQP Wang et al. (2018). MNLI, RTE, and MRPC assess the relationship between two sentences, while QNLI focuses on the relationship between a question and a sentence, and QQP evaluates the relationship between two questions.
@@ 4.2 Settings
4.3 Logical-Equivalence-Identification Contrastive Learning for Discriminative LLMs
@@ 4.4 Prompt Augmentation for Generative LLM
We assessed the robustness of AMR-LDA and LReasoner-LDA models on the PARARULE-Plus dataset Bao et al. (2022a) by modifying the test set with the contraposition law. Examples from this dataset can be found in Appendix Figures 9 and 10. AMR-LDA showed enhanced robustness on these altered tests compared to LReasoner-LDA.
@@ 8 Ethics Statement
9 Future Work
It is worth exploring how data augmentation can be used for dynamic prompt tuning in logical reasoning tasks Qi et al. (2024, 2023); Bao et al. (2020). Several studies Bao et al. (2023); Liu et al. (2023) have explored task variation formats of ReClor, LogiQA, and LogiQA-2 by altering the order of options or replacing the answers, and have found that large language models perform significantly worse under these variations. It is also worth exploring how AMR can work in conjunction with logic programming to iteratively improve reasoning performance Wang et al. (2024); Bao et al. ; Bensemann et al. (2022); Tan et al. (2023); Ni et al. (2022); Bao et al. (2022b, 2025); Bao (2025); Gendron et al. (2023). Furthermore, it is worth investigating how alternative LoRA fine-tuning methods can be used to train only the LoRA adapters Xiao et al. (2024).